# Testing the Star-disk Collision Model for Quasi-periodic Eruptions

### v1

@@ 4.2 Other QPE Sources
Next, we present results for other QPE sources. The constraints for RX J1301 are shown in the right panel of Figure 1. It shows that this source does not even have an allowed $r_{*}$-parameter space that can satisfy both Eq. (12) and Eq. (13). Considering an extremely low $\eta$ (e.g., $\sim 10^{-3}$) might allow a light-purple shaded region to appear, but it will still be above the black solid line, meaning that the star would suffer repeated partial disruptions. Therefore, the star-disk model can hardly work for RX J1301.
@@
Figure 2 shows the constraints for other four QPE sources. One can see that for eRo-QPE1, there is no parameter space that satisfies both constraints Eq. (12) and Eq. (13). For eRo-QPE2, a small parameter space of $r_{*}\sim 0.5$ for $M_{\rm{h}}\gtrsim 10^{6}\ M_{\rm{\odot}}$ is allowed by both Eq. (12) and Eq. (13), but similar to the situation for GSN 069, it is well above the solid black line, so it will suffer significant partial tidal disruptions.
@@
eRo-QPE3 and eRo-QPE4 fare much better in our test, as were shown in the lower panels of Figure 2. The best case is eRo-QPE3, whose allowed parameter space is $r_{*}\sim 1$ (i.e., a sun-like star) and $M_{\rm{h}}\lesssim 10^{7}\ M_{\rm{\odot}}$, and it is free from any tidal stripping. eRo-QPE4 also has an allowed parameter space, and taking into account the tidal stripping limit, that space shrinks to $r_{*}\sim 2$ and $M_{\rm{h}}\sim 10^{8}\ M_{\rm{\odot}}$.
4.3 QPE-like Sources with Longer Periods
@@
Now we turn to those QPE-like sources with relatively long periods: ASASSN-14ko and J0230, whose results are shown in Figure 3. Firstly, because these two sources have much longer time scales $t_{\rm{e}}$ and $P$, they require extremely large $r_{*}$ to power the luminous flares, as can be seen in Eqs. (8) and (10). In some cases, the required $r_{*}$ (i.e., $10^{3-4}$ by Eq. 13) is too large to be physically reasonable. Secondly, although a large $P$ – hence a wide orbit – puts such a star safe from a full or partial tidal disruption, there is hardly a common parameter space that can meet both Eqs. (12) and (13). Therefore, the star-disk collision model can be ruled out for these two sources. We note that a repeated partial tidal disruption model has been invoked to explain ASASSN-14ko (Payne et al., 2021, 2022, 2023; Huang et al., 2023) and J0230 (Guolo et al., 2024; Pasham et al., 2024a).
4.4 Comparison of $\ T_{\rm{p}}$
@@
Figure 4: The comparison of the peak radiation temperatures $T_{\rm{p}}$ predicted by the star-disk collision model (solid lines) and those observed (crosses with error bars) for the sample. The predicted $T_{\rm{p}}$ is calculated from Eq. (14). The horizontal error bars of the observed $T_{\rm{p}}$ correspond to the estimated $M_{\rm{h}}$ range of each individual source, respectively.
@@
We also compare the predicted peak radiation temperature $T_{\rm{p}}$ (Eq. 14) with the observed ones, which are shown in Figure 4. It is obvious that, except for ASASSN-14ko, the predicted $T_{\rm{p}}$’s for all other sources are one order of magnitude lower than the observed values. This systematic discrepancy poses a serious challenge to the star-disk collision model.
@@
Note that Eq. (14) is based on the assumption that the radiation is fully thermalized with the matter when it is diffusing out of the expanding shocked material. Linial & Metzger (2023) argue that it may not be the case, so that $T_{\rm{p}}$ could be higher.
@@
For ASASSN-14ko, the observed $kT_{\rm{p}}\sim 1\ \rm{eV}$ is low, which is not surprising since its radiation during each eruption is dominated by the optical waveband (Payne et al., 2021, 2022, 2023; Huang et al., 2023). Thus, the discrepancy does not appear to be an issue. However, we have shown in §4.3 that the star-disk collision mdoel can not pass the $r_{*}$ test for this source, thus the model can be ruled out.
5 Conclusion
The star-disk collision model proposed for QPE and QPE-like sources has attracted a lot of attention and remains a hot topic of discussion. However, the matching between the observed radiative properties, including $L_{\rm{p}}$, $t_{\rm{e}}$ and $T_{\rm{p}}$, and the model predictions has not been fully examined. In our study, we derive constraints on $r_{*}$ and $T_{\rm{p}}$ as functions of $M_{\rm{h}}$ for the star-disk collision model, using the typical values of these observables. We then apply this method to a sample of QPE and QPE-like sources, with a particular attention to what value or range of the stellar radius $r_{*}$ is required to reproduce these observables. The following are our findings.
1) Within our sample, the eruption luminosity and duration constraints require that $r_{*}\sim 3-10$ for GSN 069, $r_{*}\sim 0.3-1$ for eRo-QPE2, $r_{*}\sim 0.5-4$ for eRo-QPE3 and $r_{*}\sim 2-5$ for eRo-QPE4, respectively. There is no allowed parameter space for the rest (RX J1301, eRo-QPE1, ASASSN-14ko and J0230) of the sample.
2) Introducing the full or partial tidal disruption limit further squeezes the allowed parameter space of eRo-QPE4 to $r_{\star}\sim 2$. The best candidate source eRo-QPE3 satisfies this limit well. For GSN 069 and eRo-QPE2, however, they would have suffer significant partial disruption and the accretion of the stripped mass might be the dominant energy source powering the eruptions, thus undermining the foundation of the star-disk collision model.
3) There exists a systematic discrepancy between the model-predicted $T_{\rm{p}}$ $\sim\left(0.01\ \rm{keV}\right)$ and observed values $\sim\left(0.1\ \rm{keV}\right)$, though the incomplete thermalization of the diffusive radiation might be a remedy.
Our findings pose serious challenges to the star-disk collision model. Specifically, we find that for most of the sample, the requirements of $r_{*}$ to reproduce both $L_{p}$ and $t_{e}$, and the requirement that the star should not be partially disrupted, are in tension with each other.
@@ 5 Conclusion
One could also consider a compact object, such as a BH, as the object that collides with the disk. In this case, the eruptions will be produced by the accretion of the material that the BH gravitationally captured during the collision. The gravitational capture radius is $R_{\rm{c}}=2GM_{\rm{*}}/v_{\rm{k}}^{2}$ $\approx 5\times 10^{-4}m_{*}P_{\rm{1}}^{\frac{2}{3}}M_{\rm{6}}^{-\frac{2}{3}}\ R_{\rm{\odot}}$, and the captured mass is $M_{c}=2\pi R_{c}^{2}(\sqrt{2}\Sigma_{d})$. The total energy released from the accretion is $E_{\rm acc}=\eta M_{c}c^{2}$ $\approx 4\times 10^{39}\eta_{-1}\alpha_{-1}^{-1}\dot{m}^{-1}m_{*}^{2}P_{\rm{1}}^{\frac{7}{3}}M_{\rm{6}}^{-\frac{7}{3}}\ \rm{erg}.$
The typically observed total radiation energy during one eruption from a QPE source is $E_{\rm{rad}}\simeq 7\times 10^{45}\ \rm{erg}$ (Miniutti et al., 2023b). This would require $m_{*}\sim 10^{3}$, i.e., an intermediate mass BH, as the collider. Driven by gravitational wave radiation, the orbital decay time scale of this binary massive BH system would be $\tau_{\rm{GW}}\approx 2\times 10^{3}\left(m_{*}/10^{3}\right)^{-1}M_{\rm{6}}^{-\frac{2}{3}}P_{\rm{1}}^{\frac{8}{3}}\ \rm{yr}$ (Shapiro & Teukolsky, 1986; Peters & Mathews, 1963).
@@
This work is supported by the National Natural Science Foundation of China grant 12393814.

### v3

@@ IV.2 Other QPE Sources
Next, we present results for other QPE sources. The constraints for RX J1301 are shown in the right panel of Figure 1. It shows that this source does not even have an allowed $r_{\star}$-parameter space that can satisfy both Eq. (12) and Eq. (13). Considering an extremely low $\eta$ (e.g., $\sim 10^{-3}$) might allow a light-purple shaded region to appear, but it will still be above the black solid line, meaning that the star would suffer repeated partial disruptions. Therefore, the shock-cooling emission model from star-disk collisions can work for RX J1301.
@@
Figure 2 shows the constraints for other four QPE sources. One can see that for eRO-QPE1, there is no parameter space that satisfies both constraints Eq. (12) and Eq. (13). For eRO-QPE2, a small parameter space of $r_{\star}\sim 0.5$ for $M_{\rm{h}}\gtrsim 10^{6}\ M_{\rm{\odot}}$ is allowed by both Eq. (12) and Eq. (13), but similar to the situation for GSN 069, it is well above the solid black line, so it will suffer significant partial tidal disruptions.
@@
eRO-QPE3 and eRO-QPE4 fare much better in our test, as were shown in the lower panels of Figure 2. The best case is eRO-QPE3, whose allowed parameter space is $r_{\star}\sim 1$ (i.e., a sun-like star) and $M_{\rm{h}}\lesssim 10^{7}\ M_{\rm{\odot}}$, and it is free from any tidal stripping. eRO-QPE4 also has an allowed parameter space, and taking into account the tidal stripping limit, that space shrinks to $r_{\star}\sim 2$ and $M_{\rm{h}}\sim 10^{8}\ M_{\rm{\odot}}$.
Assuming an inflated star whose bloating factor is 33 (Yao et al., 2025), Eq. (16) becomes $r_{\rm{*}}<1.3\ n^{-\frac{9}{4}}P_{1}^{\frac{3}{2}}\left(\frac{b}{3}\right)^{-\frac{5}{4}}$. Compared with the case of no inflation $\left(b=1\right)$, its full or partial tidal disruption limit would provide a more squeezed parameter space because an inflated star would be easier to get disrupted.
IV.3 QPE-like Sources with Longer Periods
@@
Now we turn to those QPE-like sources with relatively long periods: ASASSN-14ko and J0230, whose results are shown in Figure 3. Firstly, because these two sources have much longer time scales $t_{\rm{e}}$ and $P$, they require extremely large $r_{\star}$ to power the luminous flares, as can be seen in Eqs. (8) and (10). In some cases, the required $r_{\star}$ (i.e., $10^{3-4}$ by Eq. 13) is too large to be physically reasonable. Secondly, although a large $P$ – hence a wide orbit – puts such a star safe from a full or partial tidal disruption, there is hardly a common parameter space that can meet both Eqs. (12) and (13). We note that a repeated partial tidal disruption model has been invoked to explain ASASSN-14ko (Payne et al., 2021, 2022, 2023; Huang et al., 2023).
IV.4 Comparison of $T_{\rm{p}}$
@@
Figure 4: The comparison of the peak radiation temperatures $T_{\rm{p}}$ predicted by the shock-cooling emission model from star-disk collisions (solid lines) and those observed (crosses with error bars) for the sample. The predicted $T_{\rm{p}}$ is calculated from Eq. (14). The horizontal error bars of the observed $T_{\rm{p}}$ correspond to the estimated $M_{\rm{h}}$ range of each individual source, respectively.
@@
We also compare the predicted peak radiation temperature $T_{\rm{p}}$ (Eq. 14) with the observed ones, which are shown in Figure 4. It is obvious that, except for ASASSN-14ko, the predicted $T_{\rm{p}}$’s for all other sources are one order of magnitude lower than the observed values.
@@
Note that Eq. (14) is based on the assumption that the radiation is fully thermalized with the matter when it is diffusing out of the expanding shocked material. Linial and Metzger (2023) argue that it may not be the case, so that $T_{\rm{p}}$ could be higher.
@@
For ASASSN-14ko, the observed $kT_{\rm{p}}\sim 1\ \rm{eV}$ is low, which is not surprising since its radiation during each eruption is dominated by the optical waveband (Payne et al., 2021, 2022, 2023; Huang et al., 2023). Thus, the discrepancy does not appear to be an issue. However, we have shown in §IV.3 that the shock-cooling emission model from star-disk collisions can not pass the $r_{\star}$ test for this source, thus the model can be ruled out.
V Discussion
V.1 Orbital Inclination Effect
Our above analysis so far has assumed a vertical stellar orbit, i.e., the inclination angle $i$ between the angular momenta of the star and the disk is $\frac{\pi}{2}$. Here, we consider a general case whose $i$ has other values.
In the reference frame of the star, the relative velocity of the disk is $\sqrt{2\left(1-\cos{i}\right)}v_{\rm{k}}$ and the effective length of the stellar in-disk trajectory is $\min\left[2\sqrt{\frac{2}{1+\cos{i}}}h,\ 2\pi R\right]$. For an extremely prograde orbit ($i\approx 0$), the collisions would be very weak. An intermediate inclination ($\frac{\pi}{4}\lesssim i\lesssim\frac{3\pi}{4}$) would only introduce an order-of-unity change to our previous results. Thus, for the demonstrative purpose, here we consider an extremely retrograde orbit ($i\approx\pi$) as the most optimistic case.
In this case, the relative velocity is $2v_{\rm{k}}$, the effective length of the stellar trajectory is $\approx 2\pi R$ and the cross section is $\pi R_{\star}\times\min\left[R_{\star},h\right]$. From Eq. (2), one has $h\approx 3\dot{m}M_{\rm{6}}\ R_{\rm{\odot}}$. Considering two regimes $h>R_{\star}$ and $h<R_{\star}$, respectively, one could derive the duration and the peak luminosity, similarly to in $\S$ 3.3 - 3.4, as
[…]
and
[…]
Based on Eqs. (17-18) and using Eq.(11), one can derive two solutions of $r_{\star}$ in terms of observables, similarly to Eqs. (12-13). One then checks whether the two solutions have overlapping region, considering a numerical range of $\left(0.01,0.1\right)$ for both $\alpha$ and $\eta$, and whether they satisfy the tidal constraint of Eq. (16), similarly to Figures 1-3.
We found that three sources (GSN 069, eRO-QPE2, eRO-QPE3) have allowed $r_{\star}\sim 0.1$ in the $h>R_{\star}$ regime; one source (eRO-QPE1) has allowed $r_{\star}\sim 0.01$ in the $h<R_{\star}$ regime. The rest in our sample do not have any allowed $r_{\star}$.
Therefore, the extremely retrograde orbit is indeed a feasible solution to four sources. However, one should note that for a random orientation of the stellar orbit, the differential probability distribution of $i$ is $\propto\sin i$, such that the probability for $i\in\left(\frac{3\pi}{4},\pi\right)$ is $\approx 15\%$ (also see Suzuguchi and Matsumoto, 2025). Thus, the chances for this most optimistic case are low.
V.2 Stream-disk Collisions
Here, we briefly consider the stream-disk collision, a derivative version of the original star-disk collision model, where the stream was shed from the star via ablation (Lu and Quataert, 2023; Linial and Quataert, 2024; Yao et al., 2025; Linial et al., 2025) or tidal stripping (Lu and Quataert, 2023; Huang et al., 2023).
Regarding the cross section of the stream, according to Yao et al. (2025), its radial width is determined by the dispersion in orbital energy to be $\Delta R\approx 10\ P_{\rm{1}}m_{\star}^{\frac{1}{2}}r_{\star}^{-\frac{1}{2}}\ R_{\odot}\gg R_{\star}$, whereas its transverse size is $\approx R_{\star}$ (In Linial et al. 2025, numerically calculating the Hills equations yields an even more extended stream size). Therefore, when the stream crosses the disk, it has a much larger shocking area $\pi R_{\star}\Delta R$ than the star alone.
As long as the stream’s column density is larger than the disk surface density: $\Sigma_{\rm{s}}=\frac{\Delta M_{\star}}{\pi R_{\star}\Delta R}\gtrsim\Sigma_{\rm{d}}$ (this requires $P_{\rm{1}}\lesssim 6$, while other parameters are fixed at their fiducial values; see Eq. 18 of Yao et al., 2025) where $\Delta M_{\star}$ is the stream’s mass, the stream-disk collision would shock more disk material and then power a brighter and longer-lived flare than the star-disk collision.
In this scenario, the framework in $\S$ III for constraining the shock-cooling emission still applies. To get the constraints on the stream’s cross section, we can just substitute all term of $R_{\star}^{2}$ in equations there with $R_{\star}\Delta R$, and drop off the tidal disruption constraint Eq. (16). In other words, one can interpret $r_{\star}$ in Eqs. (12-13) as the square root of this cross section in unit of $R_{\odot}$.
After removing the green and black horizontal lines in Figures 1-3, one can see that four sources within our sample (i.e., GSN 069, eRO-QPE2, eRO-QPE3 and eRO-QPE4) have the allowed cross-section size that can simultaneously reproduce the observed $L_{\rm{p}}$, $t_{\rm{e}}$ and $L_{\rm{Q}}$. The rest does not pass our tests on radiative properties. We conclude that, under the condition of $\Sigma_{\rm{s}}\gtrsim\Sigma_{\rm{d}}$, the stream-disk collision can work for the majority of the QPE sources (excluding the two long-period QPE-like events).
On the other hand, if $\Sigma_{\rm{s}}<\Sigma_{\rm{d}}$ (i.e., $P_{\rm{1}}>6$), the reverse shock propagating into the debris stream might be strong enough and the shocked stream may dominate the emission and power the eruption. In this case, the extended radial size of the stream leads to a significant dispersion in its orbital energy. Thus, different segments of the stream arrive the disk plane at markedly different times. The spread of arrival times : $\Delta t_{\rm{s}}=\left[P_{\rm{orb}}\left(R+\Delta R\right)-P_{\rm{orb}}\left(R-\Delta R\right)\right]/2\approx\frac{3}{2}\frac{\Delta R}{R}P_{\rm{orb}}\approx 6.5m_{\star}^{\frac{1}{2}}M_{\rm{6}}^{-\frac{5}{6}}\left(\frac{P_{\rm{1}}}{6}\right)^{\frac{4}{3}}\ \rm{hr}$, would represent the effective collision time scale, which might be comparable or even larger than the expansion time scale (cf. Eq. 8) of the shocked stream material.
This indicates that, for the shocked stream case, the stream-disk collision is a prolonged collision, rather different from the almost instantaneous impact in the star-disk collisions. Hence, our framework presented in $\S$ III and the findings in $\S$ IV would not apply for this case.
V.3 Break-out Emission
Besides the shock-cooling emission model, in analogy with a supernova explosion, the star-disk collision may give rise to a break-out emission (Tagawa and Haiman, 2023). The break-out emission releases the internal energy on a short time scale without the adiabatic energy loss, which is expected to produce a more luminous flare. However, the duration of break-out emission is approximated by its dynamical time scale ($\sim h/v_{\rm{k}}$), which poses a challenge in producing a QPE flare with a typical duration of $1\ \rm{hr}$. Although Tagawa and Haiman (2023) addressed the duration issue by introducing a bow shock profile, more careful checks are required in future work.
V.4 BH as the Collider
@@
Lastly, we briefly consider a compact object, such as a BH, which is free of the tidal disruption constraint, as the object that collides with the disk. In this scenario, the collider will tidally capture the disk material during each crossing. The capture radius is $R_{\rm{c}}=2GM_{\rm{*}}/v_{\rm{k}}^{2}$ $\approx 5\times 10^{-4}m_{*}P_{\rm{1}}^{\frac{2}{3}}M_{\rm{6}}^{-\frac{2}{3}}\ R_{\rm{\odot}}$ and the captured mass is $M_{\rm{c}}=2\pi R_{c}^{2}(\sqrt{2}\Sigma_{d})$.
The most efficient way to release gravitational energy is the BH accretion. Assuming the captured material is completely accreted by the collider, the total energy released is $E_{\rm acc}=\eta M_{\rm{c}}c^{2}\approx 1.5\times 10^{40}\eta_{-1}\alpha_{-1}^{-1}\dot{m}^{-1}m_{*}^{2}P_{\rm{1}}^{\frac{7}{3}}M_{\rm{6}}^{-\frac{7}{3}}\ \rm{erg}$.
The typically observed total radiation energy during one eruption from a QPE source is $E_{\rm{rad}}\simeq 7\times 10^{45}\ \rm{erg}$ (Miniutti et al., 2023b). This would require $m_{*}\sim 10^{3}$, i.e., an intermediate mass BH (IMBH), as the collider. Alternatively, if a QPE is powered by a non-accretion process, such as the shock heating and expansion-cooling of the captured material, the effective radiative efficiency $\eta$ would be lower, thus the required $m_{\star}$ would be even higher.
The gravitational wave radiation driven orbital decay time scale of this IMBH-SMBH binary would be $\tau_{\rm{GW}}\approx 2\times 10^{3}\left(m_{*}/10^{3}\right)^{-1}M_{\rm{6}}^{-\frac{2}{3}}P_{\rm{1}}^{\frac{8}{3}}\ \rm{yr}$ (Shapiro and Teukolsky, 1986; Peters and Mathews, 1963). Although this is not inconsistent with any observational fact of QPEs, the rarity of direct observational evidence of IMBHs in general so far puts this possible scenario into question. More support and independent evidence are needed for this scenario.
VI Conclusion
The star-disk collision model proposed for QPE and QPE-like sources has attracted a lot of attention and remains a hot topic of discussion. However, the matching between the observed radiative properties, including $L_{\rm{p}}$, $t_{\rm{e}}$ and $T_{\rm{p}}$, and the model predictions has not been fully examined. In our study, we derive constraints on $r_{\star}$ and $T_{\rm{p}}$ as functions of $M_{\rm{h}}$ for the shock-cooling emission model from star-disk collisions, using the typical values of these observables. We then apply this method to a sample of QPE and QPE-like sources, with a particular attention to what value or range of the stellar radius $r_{\star}$ is required to reproduce these observables. The following are our findings.
1) Within our sample, the eruption luminosity and duration constraints require that $r_{\star}\sim 3-10$ for GSN 069, $r_{\star}\sim 0.3-1$ for eRO-QPE2, $r_{\star}\sim 0.5-4$ for eRO-QPE3 and $r_{\star}\sim 2-5$ for eRO-QPE4, respectively. There is no allowed parameter space for the rest (RX J1301, eRO-QPE1, ASASSN-14ko and J0230) of the sample.
2) Introducing the full or partial tidal disruption limit further squeezes the allowed parameter space of eRO-QPE4 to $r_{\star}\sim 2$. The best candidate source eRO-QPE3 satisfies this limit well. For GSN 069 and eRO-QPE2, however, they would have suffer significant partial disruption and the accretion of the stripped mass might be the dominant energy source powering the eruptions, thus undermining the foundation of the shock-cooling emission model from star-disk collisions.
3) The influence of the orbital inclination $i$ is the most notable when the stellar orbit is extremely retrograde. In that optimistic case, a star with a small $r_{\star}$ of 0.010.10.01-0.1 would be able to reproduce the observed $t_{\rm{e}}$ and $L_{\rm{p}}$ for four sources (i.e., GSN 069, eRO-QPE1, eRO-QPE2 and eRO-QPE3). However, it is rare for the syste to have $i\sim\pi$.
4) There exists a systematic discrepancy between the model-predicted $T_{\rm{p}}$ $\sim\left(0.01\ \rm{keV}\right)$ and observed values $\sim\left(0.1\ \rm{keV}\right)$, though the incomplete thermalization of the diffusive radiation might be a remedy.
Our findings pose serious challenges to the shock-cooling emission model from star-disk collisions. Specifically, we find that for most of the sample, the requirements of $r_{\star}$ to reproduce both $L_{p}$ and $t_{e}$, and the requirement that the star should not be partially disrupted unless including the inclination effect, are in tension with each other. This issue is largely alleviated when the star is accompanied by a debris stream shed from the star and the stream-disk collisions produces the eruptions.
@@ VI Conclusion
We thank an anonymous reviewer for useful comments and suggestions that helped improve the manuscript. This work is supported by the National Natural Science Foundation of China grant 12393814.